OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (2, −1) and a directrix of y = −one half. f(x) = −(x + 2)2 − three fourths f(x) = (x + 2)2 + three fourths f(x) = −(x − 2)2 + three fourths f(x) = −(x − 2)2 − three fourths

OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (0, 1) and a directrix of y = −1. f(x) = −one fourth x2 f(x) = one fourth x2 f(x) = −4x2 f(x) = 4x2

OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (−5, 5) and a directrix of y = -1. f(x) = −one twelfth (x − 5)2 + 2 f(x) = one twelfth (x − 5)2 + 2 f(x) = −one twelfth (x + 5)2 + 2 f(x) = one twelfth (x + 5)2 + 2

OpenStudy (anonymous):

@amistre64 these are the last question that i need help on

OpenStudy (anonymous):

@AngelPSP @DeadShot @allopersonwhat @LilliBelle @nicholasradley @cupcake111 @Jonathan1997 @damien @RyGuy

OpenStudy (jdoe0001):

 recall that the directrix and the focus are both equidistant to the vertex where would that put the vertex of the parabola?

OpenStudy (anonymous):

umm i dont really know... im finna guess and say 2